# Toward Learning POMDPs Beyond Full-Rank Actions and State Observability

### v1

@@ 3.2 Transforming Predictive State Representations
Suppose we have a Hankel matrix $\mathcal{H}$, estimated in the limit of infinite data. Since the Hankel matrix $\mathcal{H}$ can be computed by multiplying two low-rank factors together, a natural first step of our method (and learning a PSR) is to compute a rank factorization of $\mathcal{H}$ (bootsClosingLearningplanning2011; balleSpectralLearning2014). One way to achieve this factorization is to compute a singular-value decomposition of the Hankel matrix $\mathcal{H}=U\Sigma V^{T}$, where singular values under a specified threshold (and their corresponding orthogonal vector components) are dropped. The truncated SVD is converted into a rank factorization by computing $A=U\Sigma$ to be the left factor and $V^{T}$ to be the right factor. Crucially, since $A\cdot V^{T}$ and $\mathbf{Forw}\cdot\mathbf{Back}$ both form rank factorizations of $\mathcal{H}$ (according to assumptions in Sec 3.3), there must exist some invertible transformation $P$ such that $A=\mathbf{Forw}\cdot P$ and $P^{-1}\cdot\mathbf{Back}=V^{T}$ (see Appendix A.2).
@@
Moving one step earlier in the Hankel construction (Sec. 3.1), we can relate transitions, observations, and initial distributions with the rank factors and the Hankel matrix using Eqs. 1-4. Let $\mathit{hists}^{ao}$ denote an ordered set of all history indices that end in action-observation pair $ao$, and $\mathit{hists}^{-ao}$ denote the same set with the same ordering but without the ending pair $ao$. From Eqs. 2 and 4, we observe for each $a\in\mathcal{A},o\in\mathcal{O}$:
@@
After applying the Moore-Penrose inverse of $A$ and $V^{T}$ to solve Eqs. 7-9, we obtain the observation-transition product, initial belief, and final summation vector up to a similarity transform. The transformed initial belief $m_{0}=b_{0}^{T}P$ is called the initial vector and the transformed summation vector $m_{\infty}=P\mathbf{1}$ is called the final vector. The product $M^{ao}=P^{-1}O^{ao}T^{a}P$ is called a linear PSR update matrix. Together, this collection of matrices and vectors forms a linear PSR model (littmanPredictiveRepresentations2001; bootsClosingLearningplanning2011). A PSR can be used to compute the likelihood of observations $o_{1},o_{2},\dots,o_{n}$ under actions $a_{1},\dots,a_{n}$ by computing the product $\operatorname{\mathrm{P}}(o_{1},\dots,o_{n}|a_{1},\dots,a_{n})=m_{0}^{T}M^{a_{1}o_{1}}\cdots M^{a_{n}o_{n}}m_{\infty}=b_{0}PP^{-1}O^{a_{1}o_{1}}T^{a_{1}}\cdot O^{a_{n}o_{n}}T^{a_{n}}P^{-1}P\mathbf{1}$, with appropriate normalizations for conditional calculations. With a few more details, the argument sketched above is a proof of a result of carlyleRealizationsStochastic1971a and later balleSpectralLearning2014. The original result given by the authors was for probabilistic automata.
Proposition 1.
[carlyleRealizationsStochastic1971a; balleSpectralLearning2014] Let $\mathcal{H}=AV^{T}$ be a rank factorization of a Hankel matrix $\mathcal{H}$ with $\operatorname{rank}(\mathcal{H})=r$ formed from a POMDP with initial state $b_{\pi}$, transition matrices $\{T^{a}\}$ and observation matrices $\{O^{ao}\}$. Suppose $m_{0},\{M^{ao},\forall(a,o)\in\mathcal{A}\times\mathcal{O}\},m_{\infty}$ are computed as in Eqs. 7, 8, and 9. Then there exists a nonsingular matrix $P\in\mathbb{R}^{r\times r}$ such that $P^{-1}M^{ao}P=O^{ao}T^{a}$ for all $a\in\mathcal{A},o\in\mathcal{O}$, $m_{0}^{T}P=b_{\pi}$, and $P^{-1}m_{\infty}=\mathbf{1}$.
3.3 Assumptions
[…]
Our assumptions have a few consequences on the estimated Hankel matrix. We only sample the starting state distribution $b_{0}$ at the start of the problem, so $b_{0}$ has little influence over the Hankel matrix. Instead, the Hankel matrix will take on the stationary distribution $b_{\pi}$ as the initial distribution instead. Furthermore, the rank of the resulting Hankel matrix will be equivalent to the number of states of the POMDPs in the restricted class that adhere to our assumptions, as opposed to the lower bound as is the case for general POMDPs. We formalize these consequences in Appendix A.
@@ 4 Computing the Similarity Transform
In many problem scenarios, $P$ is nontrivial (e.g., not the identity matrix). The rank factors computed via SVD are orthogonal matrices, which is not always the case for $\mathbf{Forw}$ and $\mathbf{Back}$. Furthermore, since rank factors may take real number values, $m_{0}$ and $\{M^{ao}\}$ are real-valued as well, and we cannot interpret their entries as state, transition, or observation emission likelihoods. These likelihoods are essential for downstream operations that direct agent behavior.
The key, then, is to recover the similarity transformation $P$. Once we know $P$, we can recover the original POMDP parameters by inverting the transform as expressed in Prop. 1. The observation and transition matrices can be recovered from products $O^{ao}T^{a}$ by computing the sums of the rows to form the diagonal of $O^{ao}$ and then normalizing the rows to form $T^{a}$. Our approach can recover $P$ up to a certain partition of states, which we introduce with an example.
A running example. Consider the POMDP illustrated in Figure 1, which is modified from the Float-Reset domain introduced by littmanPredictiveRepresentations2001. Like the original, the float action transitions the state up and down a line graph and will always emits an observation of 0. The reset action, also identical to the original, deterministically sets the state to the left end of the graph. This action emits an observation of 1 if the state is already in the leftmost state and 0 otherwise. The observations of the sense action are the same as reset, except each state of the system does not change. We also augment the system a reward function; the agent obtains +1 reward for executing any action in the state adjacent to the reset state and zero reward otherwise. This system is challenging to learn due to its nontrivial partial observability. Aside from the two leftmost states (when treating rewards as observations), all other states in this POMDP have the same observation distributions, regardless of the action. Furthermore, the transition matrix corresponding to the reset action is singular since it is zero everywhere except for a single column of ones.
@@ 4.1 Recovery up to a Full-Rank Observability Partition
Figure 1: Sense-Float-Reset. Edges are labeled with transition probabilities, and nodes are labeled with observations and reward received upon leaving the state. The reward of a state is zero unless specified otherwise. Observability partitions are represented by node shade.
Our algorithm can estimate the similarity transform $P$ up to the full-rank observability partition, which we formalize in Theorem 1. Our statement is given in the regime of infinite data; for parameters introduced for finite data, see Appendix B.1.
Theorem 1.
Let $\mathcal{H}$ be a Hankel matrix of POMDP $(\mathcal{S},\mathcal{T},\mathcal{A},\mathcal{O},\mathcal{Z},b_{\pi},R,\gamma)$ that adheres to the assumptions in Sec. 3.3, where $b_{\pi}$ is the stationary distribution under a uniform random policy $a\sim\operatorname{Unif}(\mathcal{A})$. Let $S_{\Pi}\subset 2^{\mathcal{S}}$ be the full-rank observability partition of the POMDP. Let $A$ and $V^{T}$ be a rank factorization of $\mathcal{H}$, and $m_{0}$, $\{M^{ao}:a\in\mathcal{A},o\in\mathcal{O}\}$, and $m_{\infty}$ be the linear PSR model as computed via Eqs. 7-9. Then there exists an algorithm on inputs $A$, $V^{T}$, $m_{0}$, $\{M^{ao}\}$, and $m_{\infty}$ that computes a nonsingular matrix $\tilde{P}$, such that if we let
@@
for all $a_{1},\dots,a_{n}\in\mathcal{A}$, $o_{1},\dots,o_{n}\in\mathcal{O}$, and integer $n>0$ and every partition set $S\in S_{\Pi}$.
Figure 2: An illustration of Theorem 1 applied to Sense-Float-Reset. Summing indices over partitions, represented by box shades, will compute the likelihood of the system state in that partition.
What Theorem 1 states is that we must sum over indices of the initial ‘belief vector’ to compute the likelihood the system is in a particular partition (Eq. 10). The same remains true when computing joint likelihoods between observations and the current state partition (Eq. 14; see Fig. 2 for a worked example for Sense-Float-Reset.). For POMDPs that have unique observation distributions across all actions, each state is in its own singleton partition, and we can recover the full similarity transform. Otherwise, we recover $P$ up to the full-rank observability partition. We note it is possible for us to recover some POMDPs that have fewer observations than states, since the collection of distributions over emitted observations across all actions must be distinct (see Appendix 7 for examples).
4.1.1 On the Restrictiveness of Learnable Systems
To benefit from the result of Theorem 1, the systems to be learned must satisfy the assumptions stated in Sec. 3.3 and contain full-rank actions. Here, we discuss when these assumptions are satisfied.
Full-Rank Transitions. In automated manipulation, robot actions have a desired transition state but may also fail (a gripper misses a grasp, slips of a drawer handle, etc.). One way these actions have been modeled in robot planning systems is to designate a successful ‘desired state’ with some success likelihood $p_{succ}$, and have the system state ‘fail’ with some likelihood (causing a self-transition) (kaelbling_integrated_2013; garrett_online_2020). In POMDP terms, these types of actions can be simply modeled as the convex combination $p_{\text{succ}}T+(1-p_{\text{succ}})I$, where $T$ is a matrix with rows containing all zeros except for a single entry of $1$ (the desired states), the identity $I$ indicates self-loop failure dynamics, and $p_{succ}$ the likelihood of an action succeeding. Under mild assumptions (e.g. $p_{\text{succ}}\neq 1/2,1$), these actions are full-rank (see Appendix A.6).
Ergodic Systems. Since we stipulate that POMDPs must be learned from a single trajectory, it is reasonable that the robot must be able to explore every state to correctly learn transition dynamics and observation emissions. One condition of ergodicity, irreducibility, ensures that the system does not get trapped in a subset of states. Furthermore, in many robot manipulation scenarios, robots are given a passive ‘sensing’ action that only obtains an observation sample without causing a change to the system state (kaelbling_integrated_2013). The presence of these actions break any periodic cycles, the other condition of ergodicity.
@@ 4.2 Recovering Observation Distributions from Full-Rank Actions
We now introduce an algorithm that computes the similarity transform $\tilde{P}$. Our approach is a reformulation of the tensor decomposition method (anandkumarTensorDecompositions2012; azizzadenesheliReinforcementLearning2016) for linear PSR models.
Our procedure begins by marginalizing out the observations in matrices $M^{ao}$, yielding the transitions $T^{a}$ up to similar transform $P$. This marginalization can be done by summing all matrices $M^{ao}$ over all $o\in\mathcal{O}$ for some fixed $a\in\mathcal{A}$:
@@
With a slight abuse of notation, we denote $P^{-1}T^{a}P$ as the matrix $M^{a}$. The next step of our procedure continues with transitions that are full-rank, which can easily be determined by a threshold test on the singular value decomposition on all matrices $M^{a}$. Let $\mathcal{M}_{\mathit{full}}=\{P^{-1}T^{a}P:a\in\mathcal{A}_{\mathit{full}}\}$ be the set of full-rank transitions. Next, we compute the observation matrices associated with the full-rank actions. For each $M^{a}\in M_{\mathit{full}}$ and $o\in\mathcal{O}$ we compute
@@
To reduce ambiguity, we wish to compute a joint diagonalization of all matrices $M^{ao}{M^{a}}^{-1}$, which attempts to diagonalize each matrix with the same similarity transform. We apply a method of heRandomizedMethods2024. Their method exploits the fact that sums of matrices $\{M^{ao}{M^{a}}^{-1}:a\in\mathcal{A}_{\mathit{full}},o\in\mathcal{O}\}$ do not change the invariant spaces spanned by eigenvectors of each matrix $M^{ao}{M^{a}}^{-1}$. Suppose $\{w^{ao}:a\in\mathcal{\mathcal{A}_{\mathit{full}}},o\in\mathcal{O}\}$ is a set of weights, then the weighted sum
@@
is still diagonalizable by $P$. Should we choose random weights $w_{ao}$, then the eigenvalues will be distinct up to states that share the same observation distribution almost surely. heRandomizedMethods2024 recommends sampling these weights from the unit sphere $\mathbb{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|-1}$.
Lemma 1.
@@
4.3 Recovering Partition-Level Belief State Likelihoods and Transitions
The recovered similarity transform $P^{\prime}$ formed by the eigenvectors of the random sum in Equation 18, but not the partition-level transitions. When the full-rank observability partition is nontrivial, the matrix $Q=P^{-1}P^{\prime}$ is block-diagonal, with invertible blocks that correspond to states within the same partition (see Appendix A.4 for a proof). This matrix $Q$ prevents us from using $P^{\prime}$ as the transform promised in Theorem 1. For example, when applying $P^{\prime}$ as a similarity transform to the PSR vector $m_{0}$, a restriction to the subindices of the partition $S_{1}$ yields $m_{0}P^{\prime}]_{S_{1}}=[b_{0}^{T}P^{-1}P^{\prime}]_{S_{1}}=[b_{0}^{T}]_{S_{1}}Q_{1}$, so the sum of the entries is not a proper likelihood, violating Eq. 13 of Theorem 1.
To recover partition-level likelihoods and transitions, we look to the final vector of the linear PSR after applying the transform $P^{\prime}$, e.g. $Pm_{0}=P^{\prime-1}P\mathbf{1}$. Intuitively, by applying $\operatorname{diag}(Pm_{0})$ as a similarity transform, we transform the final vector back to $\mathbf{1}$, recapturing a marginalization of the latent state variable. To avoid scenarios where $P^{\prime-1}m_{0}$ has entries of zero, we perform a pre-processing step by multiplying the system with a random block-diagonal rotation matrix $R$, whose blocks correspond to the full-rank observability partition. We take the transform $\operatorname{diag}(RP^{\prime-1}m_{\infty})RP^{\prime-1}$ as the similarity transform $\tilde{P}$ that satisfies Theorem 1 (see Appendix A.5 for proof of correctness).

### v2

@@ 3.3 Constructing a Linear Predictive State Representation
Suppose we have a Hankel matrix $\mathcal{H}$, estimated in the limit of infinite data. Since an implication of Lemma 2 is that the rank of Hankel matrix $\mathcal{H}$ is equivalent to the number of states of the POMDP, a natural first step of our method (and learning a PSR) is to compute a rank factorization of $\mathcal{H}$ [bootsClosingLearningplanning2011, balleSpectralLearning2014]. One way to achieve this factorization is to compute a singular-value decomposition of the Hankel matrix $\mathcal{H}=U\Sigma V^{T}$, where singular values under a specified threshold (and their corresponding orthogonal vector components) are dropped. The truncated SVD is converted into a rank factorization by computing $A=U\Sigma$ to be the left factor and $V^{T}$ to be the right factor. Crucially, since $A\cdot V^{T}$ and $\mathbf{Forw}\cdot\mathbf{Back}$ both form rank factorizations of $\mathcal{H}$ (according to assumptions in Sec 3.2), there must exist some invertible transformation $P$ such that $A=\mathbf{Forw}\cdot P$ and $P^{-1}\cdot\mathbf{Back}=V^{T}$ (see Appendix A.2 in the supplemental material).
@@
Moving one step earlier in the Hankel construction (Sec. 3.1), we can relate transitions, observations, and initial distributions with the rank factors and the Hankel matrix using Eqs. 1 and 2. Let $\mathit{hists}^{ao}$ denote an ordered set of all history indices that end in action-observation pair $ao$, and $\mathit{hists}^{-ao}$ denote the same set with the same ordering but without the ending pair $ao$. From Eqs. 1 and 2, we observe for each $a\in\mathcal{A},o\in\mathcal{O}$:
@@
After applying the Moore-Penrose inverse of $A$ and $V^{T}$ to solve Eqs. 5-7, we obtain the observation-transition product, initial belief, and final summation vector up to a similarity transform. The transformed initial belief $m_{0}=b_{0}^{T}P$ is called the initial vector and the transformed summation vector $m_{\infty}=P\mathbf{1}$ is called the final vector. The product $M^{ao}=P^{-1}O^{ao}T^{a}P$ is called a linear PSR update matrix. Together, this collection of matrices and vectors forms a linear PSR model [littmanPredictiveRepresentations2001, bootsClosingLearningplanning2011]. A PSR can be used to compute the likelihood of observations $o_{1},o_{2},\dots,o_{n}$ under actions $a_{1},\dots,a_{n}$ by computing the product $\operatorname{\mathrm{P}}(o_{1},\dots,o_{n}|a_{1},\dots,a_{n})=m_{0}^{T}M^{a_{1}o_{1}}\cdots M^{a_{n}o_{n}}m_{\infty}=b_{0}PP^{-1}O^{a_{1}o_{1}}T^{a_{1}}\cdot O^{a_{n}o_{n}}T^{a_{n}}P^{-1}P\mathbf{1}$, with appropriate normalizations for conditional calculations. With a few more details, the argument sketched above is a proof of a result of Carlyle and Paz [carlyleRealizationsStochastic1971a]. The original result given by the authors was for probabilistic automata, and later versions were proved for HMMs [hsuSpectralAlgorithm2012, balleSpectralLearning2014, huangMinimalRealization2016].
Proposition 1
Let $\mathcal{H}=AV^{T}$ be a rank factorization of a Hankel matrix $\mathcal{H}$ with $\operatorname{rank}(\mathcal{H})=r$ formed from a POMDP with initial state $b_{\pi}$, transition matrices $\{T^{a}\}$ and diagonal observation matrices $\{O^{ao}\}$. Suppose $m_{0},\{M^{ao},\forall(a,o)\in\mathcal{A}\times\mathcal{O}\},m_{\infty}$ are computed as in Eqs. 5, 6, and 7. Then there exists a nonsingular matrix $P\in\mathbb{R}^{r\times r}$ such that $P^{-1}M^{ao}P=O^{ao}T^{a}$ for all $a\in\mathcal{A},o\in\mathcal{O}$, $m_{0}^{T}P=b_{\pi}$, and $P^{-1}m_{\infty}=\mathbf{1}$.
@@ 4 Computing the Similarity Transform
When a reward is part of the observation, the PSR representation $M^{ao}$ can be used to compute policies that maximize the expected reward. However, if we had access to $O^{ao}$ and $T^{a}$, we could compute policies for any reward function. To recover $O^{ao}$ and $T^{a}$, we need a way to recover the similarity transform $P$ and apply Prop. 1. The observation and transition matrices can be computed from products $O^{ao}T^{a}$ by taking the sums of the rows to form the diagonal of $O^{ao}$ and then normalizing to form $T^{a}$. In many problem scenarios, $P$ is not the identity matrix, since the rank factors computed via SVD are orthogonal matrices, which is not always the case for $\mathbf{Forw}$ and $\mathbf{Back}$. Our approach can recover $P$ up to a certain partition of states, which we introduce with an example.
4.0.1 A running example.
Consider the POMDP illustrated in Figure 1, which is modified from the Float-Reset domain introduced by [littmanPredictiveRepresentations2001]. Like the original, the float action transitions the state up and down a line graph and will always emits an observation of 0. The reset action, also identical to the original, deterministically sets the state to the left end of the graph. This action emits an observation of 1 if the state is already in the leftmost state and 0 otherwise. The observations of the sense action are the same as reset, except each state of the system does not change. We also augment the system with a reward function; the agent obtains +1 reward for executing any action in the state adjacent to the reset state and zero reward otherwise. This system is challenging to learn due to its nontrivial state aliasing. Aside from the two leftmost states (when treating rewards as observations), all other states in this POMDP have the same observation distributions, regardless of the action.
@@ 4.1 Recovery up to a Full-Rank Observability Partition
Figure 1: Sense-Float-Reset. Edges are labeled with transition probabilities, and nodes are labeled with observations and reward received upon leaving the state. The reward of a state is zero unless specified otherwise. Node shade denotes observability partitions.
Our algorithm can estimate the similarity transform $P$ up to the full-rank observability partition, which we formalize in Theorem 4.1. Our statement is given in the regime of infinite data; for parameters introduced for finite data, see Appendix B.1 in the supplemental material.
Theorem 4.1
Let $\mathcal{H}$ be a Hankel matrix of POMDP $(\mathcal{S},\mathcal{T},\mathcal{A},\mathcal{O},\mathcal{Z},b_{\pi},R,\gamma)$ that adheres to the assumptions in Sec. 3.2, where $b_{\pi}$ is the stationary distribution under exploration policy $a_{t}\sim\operatorname{Unif}(\mathcal{A})$. Let $S_{\Pi}\subset 2^{\mathcal{S}}$ be the full-rank observability partition of the POMDP. Let $m_{0}$, $\{M^{ao}:a\in\mathcal{A},o\in\mathcal{O}\}$, and $m_{\infty}$ be the linear PSR model as computed via Eqs. 5-7. Then there exists an algorithm on $m_{0}$, $\{M^{ao}\}$, and $m_{\infty}$ that computes a nonsingular matrix $\tilde{P}$, such that if
@@
for all $a_{1},\dots,a_{n}\in\mathcal{A}$, $o_{1},\dots,o_{n}\in\mathcal{O}$, integer $n>0$ and partition set $S\in S_{\Pi}$.
Figure 2: An illustration of Theorem 4.1 applied to Sense-Float-Reset. Summing indices over partitions, represented by box shades, will compute the likelihood of the system state in that partition.
What Theorem 4.1 states is that we must sum over indices of the initial ‘belief vector’ to compute the likelihood the system is in a particular partition (Eq. 8). The same remains true when computing joint likelihoods between observations and the current state partition (Eq. 12; see Fig. 2 for a worked example for Sense-Float-Reset.). For POMDPs that have unique observation distributions across all actions, each state is in its own singleton partition, and we can recover the full similarity transform. Otherwise, we recover $P$ up to the full-rank observability partition. We note it is possible for us to recover some POMDPs that have fewer observations than states, since the collection of distributions over emitted observations across all actions must be distinct (see Appendix C.5.3 in the supplemental material for examples).
To benefit from the result of Theorem 4.1, the systems to be learned must satisfy the assumptions discussed in Sec. 3.2 and contain full-rank actions. Full-rank actions commonly arise in many automated manipulation contexts, our domain of interest. In automated manipulation, robot actions have a desired transition state but may also fail (a gripper misses a grasp, slips of a drawer handle, etc.). One way these actions have been modeled in robot planning systems is to designate a successful ‘desired state’ with some success likelihood $p_{succ}$, and have the system state ‘fail’ with some likelihood (causing a self-transition) [kaelbling_integrated_2013, garrett_online_2020]. In POMDP terms, these types of actions can be simply modeled as the convex combination $p_{\text{succ}}T+(1-p_{\text{succ}})I$, where $T$ is a matrix with rows containing all zeros except for a single entry of $1$ (the ‘desired states’), the identity $I$ indicates self-loop failure dynamics, and $p_{succ}$ the likelihood of an action succeeding. Under mild assumptions (e.g. $p_{\text{succ}}\neq 1/2,1$), these actions are full-rank (see Appendix A.6 in supplemental material).
@@ 4.2 Recovering Observation Distributions from Full-Rank Actions
We now introduce an algorithm that computes the similarity transform $\tilde{P}$. Our approach is a reformulation of the tensor decomposition method [anandkumarTensorDecompositions2012, azizzadenesheliReinforcementLearning2016] for linear PSR models. Our procedure begins by marginalizing out the observations in matrices $M^{ao}$, yielding the transitions $T^{a}$ up to similar transform $P$. This marginalization can be done by summing all matrices $M^{ao}$ over all $o\in\mathcal{O}$ for some fixed $a\in\mathcal{A}$:
@@
With a slight abuse of notation, we denote $P^{-1}T^{a}P$ as the matrix $M^{a}$. The next step of our procedure continues with transitions that are full-rank, which can easily be determined by a threshold test on the singular value decomposition on all matrices $M^{a}$. Let $\mathcal{M}_{\mathit{full}}=\{P^{-1}T^{a}P:a\in\mathcal{A}_{\mathit{full}}\}$ be the set of full-rank transitions. Next, we compute the diagonal observation matrices associated with the full-rank actions. For each $M^{a}\in M_{\mathit{full}}$ and $o\in\mathcal{O}$ we compute
@@
To reduce ambiguity, we wish to compute a joint diagonalization of all matrices $M^{ao}{M^{a}}^{-1}$, which attempts to diagonalize each matrix with the same similarity transform. We apply a method of He et al. [heRandomizedMethods2024]. Their method exploits the fact that sums of matrices $\{M^{ao}{M^{a}}^{-1}:a\in\mathcal{A}_{\mathit{full}},o\in\mathcal{O}\}$ do not change the invariant spaces spanned by eigenvectors of each matrix $M^{ao}{M^{a}}^{-1}$. Suppose $\{w^{ao}:a\in\mathcal{\mathcal{A}_{\mathit{full}}},o\in\mathcal{O}\}$ is a set of weights, then the weighted sum
@@
is still diagonalizable by $P$. Should we choose random weights $w_{ao}$, then the eigenvalues will be distinct up to states that share the same observation distribution almost surely. We sample these weights from the unit sphere $\mathbb{S}^{|\mathcal{A}_{\mathit{full}}|\cdot|\mathcal{O}|-1}$ [heRandomizedMethods2024].
Lemma 1
@@
4.3 Recovering Partition-Level Belief Likelihoods and Transitions
The recovered similarity transform $P^{\prime}$ formed by the eigenvectors of the random sum in Equation 16, but not the partition-level transitions. When the full-rank observability partition is nontrivial, the matrix $Q=P^{-1}P^{\prime}$ is block-diagonal, with invertible blocks that correspond to states within the same partition (see Appendix A.4 in supplemental material for a proof). This matrix $Q$ prevents us from using $P^{\prime}$ as the similarity transform promised in Theorem 4.1. For example, when applying $P^{\prime}$ as a similarity transform to the PSR vector $m_{0}$, a restriction to the subindices of the partition $S_{1}$ yields $m_{0}P^{\prime}]_{S_{1}}=[b_{0}^{T}P^{-1}P^{\prime}]_{S_{1}}=[b_{0}^{T}]_{S_{1}}Q_{1}$, so the sum of the entries is not a proper likelihood, violating Eq. 11 of Theorem 4.1.
To recover partition-level likelihoods and transitions, we look to the final vector of the linear PSR after applying the transform $P^{\prime}$, e.g. $Pm_{0}=P^{\prime-1}P\mathbf{1}$. Intuitively, by applying $\operatorname{diag}(Pm_{0})$ as a similarity transform, we transform the final vector back to $\mathbf{1}$, recapturing a marginalization of the latent state variable. To avoid scenarios where $P^{\prime-1}m_{0}$ has entries of zero, we perform a pre-processing step by multiplying the system with a random block-diagonal rotation matrix $R$, whose blocks correspond to the full-rank observability partition. We take the transform $\operatorname{diag}(RP^{\prime-1}m_{\infty})RP^{\prime-1}$ as the similarity transform $\tilde{P}$ that satisfies Theorem 4.1 (see Appendix A.5 in the supplemental material for the proof of correctness and discussion of runtime).
Summary Prop. 1 and Theorem 4.1 suggest that we can compute a PSR and then recover the corresponding system model by recovering the unknown similarity transform $P$. Sections 4.2 and 4.3 present an algorithm that applies to all POMDPs that admit the assumptions of Sec. 3.2, and computes $P$ up to the full-rank observability partition of states. The set of POMDPs learned by our method captures a large class of real-world POMDPs that arise in manipulation scenarios. It practice, it is very difficult to construct a POMDP with low-rank transitions and state aliasing and where the transitions between aliased states must be distinguished to infer a good policy.